Given that 0.475 < log base 10 of 3 < 0.478, how 
many digits are in the number 350?

Answer by AnlytcPhil(1810)

Given that 0.475 < log base 10 of 3 < 0.478, how 
many digits are in the number 350? 

We have to get 350 into the form 10x

We need to find x such that

350 = 10x

Take logs (base 10 understood) of both sides

log(350) = log(10x)

Using rules of logarithms:

50·log(3) = x

Solve for log(3)

log(3) = x/50

We are given that

0.475 < log(3) < 0.478

So substitute x/50 for log(3)

0.475 < x/50 < 0.478

Multiply thru by 50

23.75 < x < 23.9

Raise 10 to all three powers:

1023.75 < 10x < 1023.9

1023.75 < 350 < 1023.9

Integers equal to or greater than 100 but less than 101
have 1 digit.
Integers equal to or greater than 101 but less than 102
have 2 digits.
Integers equal to or greater than 102 but less than 103
have 3 digits.
Integers equal to or greater than 103 but less than 104
have 4 digits.
...      ...      ...      ...      ...      ...      ...      ...
Integers equal to or greater than 1023 but less than 1024
have 24 digits.

1023 < 1023.75 < 350 < 1023.9 < 1024

Thus 350 has 24 digits.

Checking with the calculator: 350 = 7.176879877×1023


Edwin
AnlytcPhil@aol.com